Can you please show me step by step? I am lost. I don't know where to start. I can not use excel. I have the answer. I just don't know how to get it.

a 0.05 n100 40 21 |The President of UTA is interested in estimating the proportion of Statistics students who drop BUSA 3321 this semester with an UPPER ONE-TAILED 100*(1 a)% confidence interval (a is shown on right). This semester there were n students enrolled in 3321 (shown on right) and Xo of those students dropped shown on right). Calculate the UPPER ONE-TAILED confidence limit for the true proportion who drop 3321. Round all Z scores off to 2 decimals. 0.4803 0.4960 C 0.4627 a 0.3197 0.4803 None
Answer #1

hii. i think it is not easy to explain it on excel sheet because here we can only write can not share video. but i am trying to explain deeply.

first understand how we calculate it manually.

we know the 95% one tailed confidence interval for population proportion is p+z(alpha) *s.e(p)...

now p=X0/n. calculate in excel with go to any empty cell and write =40/100, please type = sign before calculate in cell. then enter you will get 0.40

now for s.e type at any blank cell =sqrt(0.4*(1-0.4)/100) after enter you get 0.04898.. now your c. i at any blank cell type =0.4+(1.64*0.04898) you will get 0.4803.

hope you will be able to find solution. any doubt please ask freely by comment.